# A Multilingual, Culture-First Approach to Addressing Misgendering in LLM Applications

### v1

@@ Abstract
Misgendering is the act of referring to someone by a gender that does not match their chosen identity. It marginalizes and undermines a person’s sense of self, causing significant harm. English-based approaches have clear-cut approaches to avoiding misgendering, such as the use of the pronoun “they”. However, other languages pose unique challenges due to both grammatical and cultural constructs. In this work we develop methodologies to assess and mitigate misgendering across 42 languages and dialects using a participatory-design approach to design effective and appropriate guardrails across all languages. We test these guardrails in a standard large language model-based application (meeting transcript summarization), where both the data generation and the annotation steps followed a human-in-the-loop approach. We find that the proposed guardrails are very effective in reducing misgendering rates across all languages in the summaries generated, and without incurring loss of quality. Our human-in-the-loop approach demonstrates a method to feasibly scale inclusive and responsible AI-based solutions across multiple languages and cultures.
@@ 1 Introduction
In this study, we address the challenge of minimizing misgendering across 42 languages and dialects within an AI-based application. The full set of languages can be found in Appendix A. To tackle this problem, we begin by establishing guardrails in consultation with native speakers, and test them in a meeting transcript summarization context. Specifically, given a meeting and participant information, the task is to generate an accurate summary with an LLM that respects the gender identities of all participants. [^2] Data, rubrics, and prompts may be found in https://anonymized/url
@@ 1.1 Contributions
    A set of guardrails developed following participatory design practices across 42 languages and dialects, designed to avoid and mitigate misgendering rates in LLM-based systems.
@@
As part of our analysis, we also show that LLMs do not always align with humans in parts of this task, particularly around detecting misgendering. Our guardrail development likewise notes that solving this problem is complex, but not infeasible. Even though addressing misgendering over gendered or formally regulated languages goes beyond linguistics, participatory design informs deeper understanding of local culture and customs, leading to better system performance. We note that our methodologies are also extensible to other languages and contexts.
@@ 3 Guardrail Creation
For the mining step, we resourced a Writing Style Guide with instructions for writing inclusive documents in each language under consideration. This guide is used as part of the application’s internationalization efforts. [^3] https://anonymized/url
@@ 4.1 Meeting Transcript Data Generation
The data generator takes in one of 20 hypothetical meeting scenarios (a topic and a list of participants) and a language, and generates a corresponding transcript in said language. The genders of the participants are either specified as male, female, non-binary or genderfluid; or explicitly stated to be unspecified in the scenario description. The specifications serve as ground truth for verification and evaluation; and at least one of the participants was requested to be of a gender that could not readily be identified from the transcript in some scenarios. We used the same 20 scenarios for all languages. An example scenario and participant list would be as follows: Generate a dialogue in LANG between three named doctors, two of whom are male and one is female. They are discussing administrative changes coming into effect at the hospital starting next week.
@@
This pipeline is iterative: if the verifier flags the transcript as correct, we keep it. Otherwise, call the generator again with the same parameters, and, to ensure that it does not make the same mistakes, we append verifier feedback in the generator prompt. We iterate 3-4 times per language until most of the transcripts are verified to be accurate by the verifier, and obtain 20 verified transcripts per language.
Next, native speakers generate the gold transcripts by verifying and correcting if needed the verified transcripts. In this step they are requested to edit the transcript to ensure consistent gendering, paying special attention to the participant whose gender could not be readily identified. In situations where this editing could lead to loss of meaning, we consider it acceptable. Details on the annotation practices are in Section 8.
Our final dataset is 20 gold transcripts per language, which we use for our experiments going forward.
@@ 5.1 Guardrail Effectiveness
5.2 Ablation by Language Availability
@@
We compared the responses of the individual GPT-4o evaluator to human evaluation results, by ablating by the language availability classes defined by Joshi et al. (2020), shown in Figures 4, 5 and 6. The authors define a scale between 0 (lowest) and 5 (highest) to determine the level of resources of a given language by taking into account both labeled and unlabeled data available for that language. The languages with their classes are in Appendix A. We found that, although there was a clear disparity in the performance of GPT-4o when compared to humans, especially in guardrail-enabled transcripts, this was not strictly tied to a language class, and rather to the metric. Nonetheless, it was clear that the model had trouble with class 1 languages, such as Welsh, often overshooting its responses. Disparity across multiple languages was more clear in the GM score, and remained steady in the Q score.
@@ 5.3 Discussion
Overall, the language-specific guardrails created in consultation with native speakers were effective at reducing misgendering across all languages without impacting the quality of the summaries. In some cases, they slightly increased it. There were more consistent reductions in GA scores versus GM scores in both human and LLM-evaluations. This indicates that the guardrails should be adjusted to reduce gender mistakes, especially for those languages and metrics where the LLM or human evaluation show regressions. While our guardrails led to improvements in gender assumptions and errors, they could be further reduced. Additionally, incorporating more aspects of gender and misgendering can enhance the effectiveness of the guardrails.
In terms of agreement between human and LLM-evaluators, as noted from previous research, LLMs did not always align with humans, especially in some languages and metrics. This disparity was evident in low-resource languages, like Welsh, where the model has a much higher misgendering detection rate. We also observed that the SBS LLM-evaluation for the Q metric suffered from position bias. While instruction-tuning on evaluation tasks may enhance LLM-evaluators, it is advised not to depend exclusively on LLM-based evaluators in multilingual settings and nuanced cases to obtain an accurate assessment of the efficacy of solutions.
Future directions include enhancing guardrails by incorporating human evaluator feedback; and improving LLM-evaluators along with other components of the automated pipeline to increase efficiency. This approach aims to scale across multiple languages and cultures while maintaining human involvement within a feasible budget.
@@ 6 Conclusion
The act of misgendering is an exclusionary practice that could have severe impact on the people affected. It is, however, a complex subject to address in a multilingual and multicultural setting. This is not just a linguistic issue where grammatical constructions may perpetuate these behaviors, but also a sociopolitical matter, as certain mechanisms could encourage erasure. For example, imposing two genders implies that those who identify with neither are forced to choose one.
In this paper we created a process by which to address this problem while retaining cultural sensitivities. We found that our approach is effective at both reducing misgendering and retaining user satisfaction as measured by the quality of the generated text. We also found that LLMs are not always good at detecting misgendering, and this may extend to other nuanced judgments that are language and culture-specific. While this could make them unreliable detectors, for generation this means that they must be tuned from the start by incorporating appropriate instructions in the prompt.
While our work is narrowed down to a single application, in a broader context recall that any natural language is in constant flux, and change is a natural part of its development. Resistance to this change could come from the regulatory bodies themselves–consider, for example, how in Spanish the RAE mentioned that andro-centrism in the language was subjective and that using the masculine as the default collective, mixed-gender noun did not imply a loss of information on the referents Real Academia Española (2020). Detractors could be other sectors of the population: in Swedish [^4] Swedish is no longer gendered since it has two grammatical genders (common and neuter); however, it has pronouns based on natural gender (han/hon). the gender-neutral pronoun “hen”, proposed in 1966, was banned by some media outlets as early as 2012 Gustavsson (2012).
@@ 7 Limitations
Although our work is comprehensive in the sense that it is carefully constructed end-to-end, employing native speakers from the inception of the work to two (transcript and annotation) verification steps, this comes at a downside: the dataset per language is comparatively small. We have addressed this by employing measures of statistical significance whenever possible, and, as mentioned, qualitative analyses (typical for low-volume work) or participatory design on some or all steps of the process. Although we employed a single LLM for our evaluation, it must be noted that the focus of our work is less around the capabilities of LLMs themselves as data annotators/verifiers, and more on demonstrating that reducing multilingual misgendering rates is achievable.
The second limitation is the number of languages and dialects covered. While comprehensive, it does exclude large numbers of families and dialects: a glaring omission would be, for example, the Bantu languages, such as Kiswahili, which have up to 18 noun classes. We leave this for (highly encouraged) future work.
@@ 8 Ethics
All aspects of our work were reviewed and approved by an Institutional Review Board at our organization. The survey and annotation work were carried out by an external agency that recruited the participants. While the survey work did not explicitly require professional annotators, the transcript verification/correction and labelling did. All participants were compensated for their work, with salaries depending on location and seniority, starting at \$22 USD/hour. All data, including qualitative feedback and demographics, have been anonymized and will be released under a permissive license.

### v2

@@ Abstract
Misgendering is the act of referring to someone by a gender that does not match their chosen identity. It marginalizes and undermines a person’s sense of self, causing significant harm. English-based approaches have clear-cut approaches to avoiding misgendering, such as the use of the pronoun “they”. However, other languages pose unique challenges due to both grammatical and cultural constructs. In this work we develop methodologies to assess and mitigate misgendering across 42 languages and dialects using a participatory-design approach to design effective and appropriate guardrails across all languages. We test these guardrails in a standard LLM-based application (meeting transcript summarization), where both the data generation and the annotation steps followed a human-in-the-loop approach. We find that the proposed guardrails are very effective in reducing misgendering rates across all languages in the summaries generated, and without incurring loss of quality. Our human-in-the-loop approach demonstrates a method to feasibly scale inclusive and responsible AI-based solutions across multiple languages and cultures. We release the guardrails and synthetic dataset encompassing 42 languages, along with human and LLM-judge evaluations, to encourage further research on this subject.
@@ 1 Introduction
In this study, we address the challenge of minimizing misgendering across 42 languages and dialects within an AI-based application. The full set of languages can be found in Appendix A. To tackle this problem, we begin by establishing guardrails in consultation with native speakers, and test them in a meeting transcript summarization context.  [^2] Guardrails, synthetic data and evaluation data can be found at https://github.com/microsoft/Multilingual-Culture-First-Misgendering-Guardrails
@@ 1.1 Contributions
    A set of guardrails developed following participatory design practices across 42 languages and dialects, designed to avoid and mitigate misgendering rates in LLM-based systems. We release these guardrails under a permissive license for the community to use in LLM-based applications.
@@
As part of our analysis, we also show that LLMs do not always align with humans in parts of this task, particularly around detecting misgendering. Our work targets a specific LLM-based application, but the guardrails we create are generic and can be included in other applications. Our work shows that solving this problem is complex, but not infeasible. Addressing misgendering over gendered or formally regulated languages goes beyond linguistics; participatory design informs deeper understanding of local culture and customs, leading to better system performance. Our methodologies are adaptable to other languages and contexts, providing a general framework for incorporating community input into Responsible AI.
@@ 3 Guardrail Creation
For the mining step, we resourced a Writing Style Guide with instructions for writing inclusive documents in each language under consideration. This guide is used as part of the application’s internationalization efforts. [^3] https://learn.microsoft.com/en-us/style-guide
@@ 4.1 Meeting Transcript Data Generation
The data generator takes in one of 20 hypothetical meeting scenarios (a topic and a list of participants) and a language, and generates a corresponding transcript in said language. The meeting scenarios consist of conversations between participants in various settings, such as hospitals, schools, workplaces and between friends and were created by the authors of this work. The genders of the participants are either specified as male, female, non-binary or genderfluid; or explicitly stated to be unspecified in the scenario description. An example of a scenario is as follows: Generate a dialogue in <language> between three named people: a man, a woman, and someone whose gender cannot readily be identified. They are discussing the launch of a new soap brand and the projected revenue. The gender specifications serve as ground truth for verification and evaluation; and at least one of the participants was requested to be of a gender that could not readily be identified from the transcript in some scenarios. We used the same 20 scenarios for all languages.
@@
This pipeline is iterative: if the verifier flags the transcript as correct, we keep it. Otherwise, call the generator again with the same parameters, and, to ensure that it does not make the same mistakes, we append verifier feedback in the generator prompt. We iterate 3-4 times per language until most of the transcripts are verified to be accurate by the verifier, and obtain 20 transcripts per language.
Next, native speakers generate the gold transcripts by verifying and correcting if needed the final transcripts from the iterative process. In this step they are requested to edit the transcript to ensure consistent gendering, paying special attention to the participant whose gender could not be readily identified. Details on the annotation practices are in Section 8. Our final dataset is 20 gold transcripts per language, which we use for our experiments going forward.
@@ 5.1 Guardrail Effectiveness
5.2 Ablation by Language Resources
@@
We compared the responses of the individual GPT-4o evaluator to human evaluation results, by ablating by the language resource classes defined by Joshi et al. (2020), shown in Figures 4, 5 and 6. The authors define a scale between 0 (lowest) and 5 (highest) to determine the level of resources of a given language by taking into account both labeled and unlabeled data available for that language. The languages with their classes are in Appendix A. We found that, although there was a clear disparity in the performance of GPT-4o when compared to humans, especially in guardrail-enabled transcripts, this was not strictly tied to a language class, and rather to the metric. Nonetheless, it was clear that the model had trouble with class 1 languages, such as Welsh, often overshooting its responses. Disparity across multiple languages was more clear in the GM score, and remained steady in the Q score.
@@ 5.3 Discussion
Overall, the language-specific guardrails created in consultation with native speakers were effective at reducing misgendering across most languages without impacting the quality of the summaries. There were more consistent reductions in GA scores versus GM scores in both human and LLM-evaluations. This indicates that the guardrails should be adjusted to reduce gender mistakes, especially for those languages and metrics where the LLM or human evaluation show regressions. While our guardrails led to improvements in gender assumptions and errors, they could be further reduced. Additionally, incorporating more aspects of gender and misgendering can enhance the effectiveness of the guardrails.
LLM-judges did not always align with humans, especially in some languages and metrics. This disparity was evident in low-resource languages, like Welsh, where the model has a much higher misgendering detection rate. We also observed that the SBS LLM-evaluation for the Q metric suffered from position bias. While instruction-tuning on evaluation tasks may enhance LLM-evaluators, it is advised not to depend exclusively on LLM-based evaluators in multilingual settings and nuanced cases to obtain an accurate assessment of the efficacy of solutions.
Future work includes expanding the evaluation to more models, enhancing guardrails by incorporating human feedback, evaluating the effectiveness of the guardrails across other applications and improving LLM-evaluators along with other components of the automated pipeline to increase efficiency. This approach aims to scale across multiple languages and cultures while maintaining human involvement within a feasible budget.
@@ 6 Conclusion
The act of misgendering is an exclusionary practice that could have severe impact on the people affected. It is, however, a complex subject to address in a multilingual and multicultural setting. In this paper we created a process by which to address this problem while retaining cultural sensitivities. We found that our approach is effective at both reducing misgendering and retaining user satisfaction as measured by the quality of the generated text. We also found that LLMs are not always good at detecting misgendering, and this may extend to other nuanced judgments that are language and culture-specific. While this could make them unreliable detectors, for generation this means that they must be tuned from the start by incorporating appropriate instructions in the prompt.
Natural language is in constant flux, and change is a natural part of its development. Resistance to this change could come from the regulatory bodies themselves–consider, for example, how in Spanish the RAE mentioned that andro-centrism in the language was subjective and that using the masculine as the default collective, mixed-gender noun did not imply a loss of information on the referents Real Academia Española (2020). Detractors could be other sectors of the population: in Swedish [^4] Swedish is no longer gendered since it has two grammatical genders (common and neuter); however, it has pronouns based on natural gender (han/hon). the gender-neutral pronoun “hen”, proposed in 1966, was banned by some media outlets as early as 2012 Gustavsson (2012).
@@ 7 Limitations
Although our work is comprehensive in the sense that it is carefully constructed end-to-end, employing native speakers from the inception of the work to two (transcript and annotation) verification steps, this comes at a downside: the dataset per language is comparatively small. We have addressed this by employing measures of statistical significance whenever possible, and, as mentioned, qualitative analyses (typical for low-volume work) or participatory design on some or all steps of the process. Although we employed a single LLM for our evaluation, it must be noted that the focus of our work is less around the capabilities of LLMs themselves as data annotators/verifiers, and more on demonstrating that reducing multilingual misgendering rates is achievable. Similarly, although we focus on a single application where misgendering causes harm, the guardrails we produce are generic and can be used across differet applications.
Another limitation of our work is the number of languages and dialects covered. While comprehensive, it does exclude large numbers of families and dialects: a glaring omission would be, for example, the Bantu languages, such as Kiswahili, which have up to 18 noun classes. We leave this for (highly encouraged) future work.
Potential risks: Although we design the guardrails and experiments carefully, it is possible that guardrails may increase misgendering or have other unwanted effects when deployed in LLM applications. We recommend testing the guardrails with appropriate metrics for misgendering and quality before wide-scale deployment.
@@ 8 Ethics
All aspects of our work were reviewed and approved by an Institutional Review Board at our organization. The survey and annotation work were carried out by an external agency that recruited the participants. While the survey work did not explicitly require professional annotators, the transcript verification/correction and labelling did. All participants were compensated for their work, with salaries depending on location and seniority, starting at \$22 USD/hour. Out of the participants, 60% of them self-identified240 as women; 2% as non-binary; 76% had linguistics training; and 50% were between 30-45 years old. All participants were native speakers of the languages they annotated.
Dataset: All data, including qualitative feedback, has been anonymized and has been released under a permissive license. The guardrails are intended to be used within prompts of LLM applications where misgendering is a concern. We recommend using language-specific guardrails for each language separately, rather than using all guardrails in a single prompt. We recommend testing the guardrails within the application under consideration by using appropriate metrics for misgendering and quality before wide-scale deployment. The meeting transcript dataset can be used to replicate our experiment and is not intended to be used for other purposes. The human and LLM-judge data was collected to compare agreement and while it can be potentially used to fine-tune LLM-judges, we have not tested its effectiveness in improving LLM judges. The scope of the guardrails are limited to reducing misgendering as defined by the paper, and we do not address other, more general Responsible AI or bias concerns in the guardrails.